# ExploRLLM: Guiding Exploration in Reinforcement Learning with Large Language Models

### v3

@@ Abstract
In robot manipulation tasks with large observation and action spaces, reinforcement learning (RL) often suffers from low sample efficiency and uncertain convergence. As an alternative, foundation models have shown promise in zero-shot and few-shot applications. However, these models can be unreliable due to their limited reasoning and challenges in understanding physical and spatial contexts. This paper introduces ExploRLLM, a method that combines the commonsense reasoning of foundation models with the experiential learning capabilities of RL. We leverage the strengths of both paradigms by using foundation models to obtain a base policy, an efficient representation, and an exploration policy. A residual RL agent learns when and how to deviate from the base policy while its exploration is guided by the exploration policy. In table-top manipulation experiments, we demonstrate that ExploRLLM outperforms both baseline foundation model policies and baseline RL policies. Additionally, we show that this policy can be transferred to the real world without further training. Supplementary material is available at https://explorllm.github.io.
@@ I INTRODUCTION
Foundation models (FMs) [1], which refer to models trained on large-scale data (e.g., Large Language Models or Vision-Language Models), have shown significant promise in robotics. Large Language Models (LLMs), such as GPT-4 [2], can generate commonsense-aware reasoning in various scenarios. For instance, LLMs have demonstrated zero-shot planning capabilities [3], breaking down complex tasks into detailed step-by-step plans without additional training. When integrated with Vision-Language Models (VLMs), LLMs leverage cross-domain knowledge for robot perception and planning in manipulation tasks [4]. This synergy allows for the extraction of environmental affordances and constraints, forming a foundation for subsequent robotic planning [5]. Despite the impressive results of FMs, unpredictable failures in LLM predictions can still lead to robotic errors, and LLMs generally do not learn from past experiences [6, 7].
@@
Reinforcement Learning (RL) offers a powerful framework for learning decision-making and control policies through interaction with the environment [8]. However, RL struggles with the “curse of dimensionality,” where large observation and action spaces slow exploration and convergence. To address this, we propose combining FMs and RL, using FMs to guide the RL agent’s exploration. While actions generated by FMs may be sub-optimal or fail, they highlight meaningful regions in the action space for exploration. Traditional RL exploration strategies (e.g., $\epsilon$-greedy, Boltzmann exploration [9]) are stochastic, focusing on exploration-exploitation trade-offs, but lack mechanisms to incorporate prior knowledge for faster convergence. We instead use LLMs as few-shot planners, generating actions that serve as exploration steps in RL, increasing the likelihood of successful states and gathering more relevant state-action pairs for off-policy RL agents.
Our method, ExploRLLM, improves performance by compensating for FMs’ sub-optimality and biases through RL, while FMs accelerate RL training by reducing observation spaces and guiding exploration. To summarize, our main contributions are: 1) We propose ExploRLLM, which employs an RL agent with a) residual action and observation spaces based on affordances identified by FMs and b) LLM-guided exploration. 2) We introduce a prompting method for LLM-based exploration using hierarchical language-model programs, leading to faster convergence. 3) We show that ExploRLLM outperforms policies derived solely from LLMs and VLMs and generalizes to unseen scenarios, tasks, and real-world settings without additional training.
@@ II-A Foundation Models for Planning in Robotics
Researchers have shown that LLMs can generate zero-shot or few-shot plans using reasoning capabilities [3, 10], which is crucial for high-level planning in robotics. These models facilitate task-level planning by integrating environmental groundings, such as affordance value scores [11] or feedback [12], with their language groundings. Furthermore, LLMs can generate robot-centric code programs as representations for both task-level [13] and skill-level planning [14]. VLMs are increasingly integrated into robotics as a perception module of environmental context. The integration of knowledge from LLMs and VLMs facilitates the creation of perception-planning pipelines [4] and the construction of 3D value maps for zero-shot planning frameworks [5]. However, directly applying VLMs and LLMs to zero-shot tasks may not guarantee success or safety due to real-world uncertainty. In our research, we treat these actions as exploratory behaviors within an RL framework.
@@ II-B Foundation Models and Reinforcement Learning
Incorporating FMs into RL frameworks has notably improved RL’s effectiveness. In  [15], the authors have implemented LLMs as proxy reward functions, demonstrating their utility in RL. In the context of RL for robotics, LLMs are also capable of generating reward signals for robot actions by connecting commonsense reasoning with low-level actions [16], self-refinement [17] and evolutionary optimization over reward code to enable complex tasks such as dexterous manipulation [18]. Regarding exploration, authors in [19] reward RL agents toward human-meaningful intermediate behaviors by prompting an LLM. LLMs are also utilized as an intrinsic reward generator to guide exploration for long horizon manipulation tasks [20]. Contrary to these studies, our approach employs LLM-generated code policies to guide exploratory actions rather than focusing on reward shaping.
@@ III Problem Formulation
In this study, we focus on tabletop manipulation tasks to implement and evaluate our ExplorLLM method, as these tasks facilitate the creation of a base policy using FMs. We use a VLM for object detection and an LLM to identify the object to be manipulated from a command. Each manipulation task begins with a linguistically described goal, denoted by $\bm{l}$. At each time step $t$, the agent receives an observation $\bm{o}^{t}$, consisting of an overhead RGB-D image and the state of the end-effector. Similar to existing methods (e.g., Transporter [21]), the action space involves a pick and a place primitive, denoted as $\{\mathcal{P}_{\mathrm{pick}},\mathcal{P}_{\mathrm{place}}\}$, with each action parameterized by pick and place positions in a top-down view. We simplify this to a single motion primitive—either pick or place. This simplification makes the RL challenge more tractable by eliminating the need to learn a feature representation for each primitive individually. The pick or place action is defined as a tuple containing the primitive index $k$ (0 for pick, 1 for place) and a top-down view position, expressed as $\bm{x}$, i.e., $\bm{a}^{t}=(k^{t},\bm{x}^{t})$. At each time step, the agent receives a reward $r$ consisting of a dense reward component $r^{d}$ and a sparse reward $r^{s}$.
@@ IV-A Observation and Action Spaces
Our methodology leverages the strengths of LLMs and VLMs to extract the observation space used for the RL framework, as depicted in Figure 2. LLMs reformulate user-provided language commands into predefined templates and highlight the objects within these templates to form an interpreted command vector $\tilde{\bm{l}}$. For example, it identifies the pick object in a template like “Put [pick_object] on the [place_object].” It is important to note that, within a given task setting, the number and category of objects do not change. Utilizing VLMs as open-vocabulary object detectors, our system identifies and encloses objects relevant to the task within bounding boxes from the image in raw observation space $\bm{o}^{t}$, represented by their locations $\bm{X}_{\mathrm{vlm}}=[\bm{x}_{\mathrm{vlm_{1}}},\bm{x}_{\mathrm{vlm_{2}}},...]$. RGB-D visual inputs are segmented into crops based on bounding box positions, denoted as $\bm{M}_{\mathrm{vlm}}=[\bm{m}_{\mathrm{vlm_{1}}},\bm{m}_{\mathrm{vlm_{2}}},...]$. This method improves the system’s robustness to detection inaccuracies and varying object shapes. The interpreted commands $\tilde{\bm{l}}$, the positional data $\bm{X}_{\mathrm{vlm}}$ and the image patches $\bm{M}_{\mathrm{vlm}}$ are then integrated into the reformulated RL observation $\bm{s}^{t}$.
As the VLM already extracts each object’s position $\bm{X}_{\mathrm{vlm}}^{t}[i^{t}]$, the action space is converted into an object-centric residual action space, as shown in Figure 2. The reformulated action space consists of a primitive index $k$, an object index $i$ and a residual position $\bm{x}_{\mathrm{res}}$, expressed as $\tilde{\bm{a}}^{t}=(k^{t},i^{t},\bm{x}_{\mathrm{res}}^{t})$, where $\bm{x}^{t}=\bm{X}_{\mathrm{vlm}}^{t}[i^{t}]+\bm{x}_{\mathrm{res}}^{t}$. This residual action allows to pick or place objects at specific locations. This is, for example, needed when picking the letter O, and $\bm{X}_{\mathrm{vlm}}^{t}[i^{t}]$ denotes the center of the bounding box. In this case, the residual action $\bm{x}_{\mathrm{res}}^{t}$ is needed to prevent picking the letter O at its empty center.
Algorithm 1 Exploration strategy $\pi^{\text{E}}$
Input: state $\bm{s}^{t}$, LLM policies $\pi_{H}^{\text{LLM}}$, $\pi_{L}^{\text{LLM}}$
Parameter: threshold $\epsilon$
Output: action $\tilde{\bm{a}}^{t}$
1:  Sample a random number $j$ from $U(0,1)$
2:  if $j\leq\epsilon$ then
3:     Run LLM-generated high level action policy $\pi_{H}^{\text{LLM}}$$\bm{a}_{H}^{t}=(k^{t},i^{t})=\pi_{H}^{\text{LLM}}(\bm{s}^{t})$
4:     Run LLM-generated low level action policy $\pi_{L}^{\text{LLM}}$$\bm{x}_{\mathrm{res}}^{t}=\pi_{L}^{\text{LLM}}(\bm{s}^{t},\bm{a}_{H}^{t})$ $\tilde{\bm{a}}^{t}=(k^{t},i^{t},\bm{x}_{\mathrm{res}}^{t})$
5:  else
6:     Run the reinforcement learning policy $\pi^{\text{RL}}$$\tilde{\bm{a}}^{t}=\pi^{\text{RL}}(\bm{s}^{t})$
7:  end if
8:  return action $\tilde{\bm{a}}^{t}$
@@ IV-B LLM-Based Exploration
Traditional deep RL algorithms (e.g., SAC [22], PPO [23]) do not inherently promote frequent visits to high-value states in high-dimensional state-action spaces, making vision-based tabletop manipulation tasks particularly challenging. In such cases, RL agents may struggle when successful outcomes are rare. Leveraging the planning capabilities of LLMs and the perception strengths of VLMs can help guide the exploration process more effectively by tapping into the rich prior knowledge within these FMs. The LLM-based exploration strategy, denoted as $\pi^{\text{E}}$ in Algorithm 1, draws inspiration from the $\epsilon$-greedy strategy. Specifically, during the rollout collection at each timestep, the off-policy RL agent employs the LLM-based exploration technique if a sampled random variable falls below the threshold $\epsilon$. Otherwise, the action is selected according to the current RL agent’s policy, $\pi^{\text{RL}}$, as detailed in Algorithm 1.
Prior research frequently prompts LLMs at every step to create plans for robotic manipulation, making the process highly resource-intensive. This method incurs significant time and financial costs due to the numerous LLM invocations required to train a single RL agent. Drawing inspiration from CaP [14], our methodology employs the LLM to hierarchically generate language model programs, which are then executed iteratively during the training phase as exploratory actions. The hierarchical language model programs include both high-level $\pi_{H}^{\text{LLM}}$ and low-level $\pi_{L}^{\text{LLM}}$ policy code programs. A high-level plan primarily involves selecting robot action primitives and the objects to interact with based on the current state of the robot and the objects.
In contrast to high-level tasks, instructing low-level actions poses a more significant challenge because high-level states and actions are more accessible and can be represented as language. When dealing with low-level actions, the complexity of the state becomes considerably more intricate, particularly for image-based problems. Therefore, instead of a deterministic code policy, we instruct the LLM to produce a code policy $\pi_{L}^{\text{LLM}}$ for generating an affordance map according to the input image. The low-level exploration behavior is derived from a stochastic policy that relies on the values within this affordance map. Although the code generated by LLMs lacks guaranteed feasibility and accuracy in robot environments, these models can generate potentially useful policy candidates, with the one exhibiting the highest success rate being selected as shown in Figure 3.
@@ V Implementation
The main components for the implementation of ExploRLLM are RL agent, VLM-based object detection, and code policy generation by LLM.
V-1 Reinforcement learning agent
We use the Soft Actor-Critic (SAC) algorithms with modifications in the collecting rollout phase, detailed in Algorithm 1. Other implementation aspects remain consistent with the standard SAC approach in stable-baselines3 [24]. We employ two convolutional layers to transform every image patch into a vector $\bm{\phi}\in\mathbb{R}^{n\times d}$, where n is the number of objects captured by VLM and $d$ the dimension of each patch as encoded by the CNN. The vector is subsequently concatenated with the position, robot gripper state, and the extracted episodic language goal $\bm{\tilde{l}}$ to form a new vector $\bm{\phi}^{\prime}\in\mathbb{R}^{n\times d^{\prime}}$, where $d^{\prime}$ denotes the dimension of each patch’s vector following encoding and concatenation. It then goes to a self-attention layer. The output features from this layer then go into a two-layer MLP. The aforementioned structure is consistently utilized across all actor and critic networks.
V-2 VLM detection
Utilizing an open-vocabulary object detector ViLD [25], objects in the environment can be identified by given specific labels. However, implementing this model online during training is time-consuming, so ViLD is utilized solely in the evaluation phase. In the training phase, the ground truth in the simulation is used to determine the center positions of the bounding boxes. It is important to note that ViLD’s position detection in real-world scenarios is not always flawless. To simulate this imperfection, noise following a Gaussian distribution with a standard deviation equivalent to half the radius of the image crop is applied to the ground truth positions.
V-3 Code policy generation by LLM

### v4

@@ Abstract
In robot manipulation, Reinforcement Learning (RL) often suffers from low sample efficiency and uncertain convergence, especially in large observation and action spaces. Foundation Models (FMs) offer an alternative, demonstrating promise in zero-shot and few-shot settings. However, they can be unreliable due to limited physical and spatial understanding. We introduce ExploRLLM, a method that combines the strengths of both paradigms. In our approach, FMs improve RL convergence by generating policy code and efficient representations, while a residual RL agent compensates for the FMs’ limited physical understanding. We show that ExploRLLM outperforms both policies derived from FMs and RL baselines in table-top manipulation tasks. Additionally, real-world experiments show that the policies exhibit promising zero-shot sim-to-real transfer. Supplementary material is available at https://explorllm.github.io.
@@ I Introduction
Foundation Models (FMs) [1], which refer to models trained on large-scale data, have shown great potential in robotics. In particular, language-based FMs, such as Large Language Models (LLMs) and Vision-Language Models (VLMs), are increasingly used in the field. Large Language Models, such as GPT-4 [2], can generate commonsense-aware reasoning in various scenarios. For instance, LLMs have demonstrated zero-shot planning capabilities [3], breaking down complex tasks into detailed step-by-step plans without additional training. When integrated with VLMs, LLMs leverage cross-domain knowledge for robot perception and planning in manipulation tasks [4]. This synergy allows for extracting environmental affordances and constraints, forming a foundation for subsequent robotic planning [5]. Despite the impressive results of FMs, unpredictable failures in LLM predictions can still lead to robotic errors, and LLMs generally do not learn from past experiences [6, 7].
@@
On the other hand, Reinforcement Learning (RL) offers a powerful framework for learning decision-making and control policies through interaction with the environment [8]. However, RL struggles with the “curse of dimensionality,” where large observation and action spaces slow down exploration and convergence. To address this, we propose combining FMs and RL by using FMs to guide the RL agent’s exploration as depicted in Figure 1. While actions generated by FMs may be suboptimal or fail, they can highlight meaningful regions in the action space for exploration. Traditional RL exploration strategies (e.g., $\epsilon$-greedy, Boltzmann exploration [9]) are stochastic, focusing on exploration-exploitation trade-offs, but lack mechanisms to incorporate prior knowledge for faster convergence. Instead, we use LLMs as few-shot planners, generating actions that serve as exploration steps in RL, increasing the likelihood of successful states and gathering more relevant state-action pairs for off-policy RL agents.
Our method, ExploRLLM, improves performance by compensating for FMs’ sub-optimality and biases through RL, while FMs accelerate RL training by reducing observation spaces and guiding exploration. To summarize, our main contributions are the following.
  1. 1.
    We propose ExploRLLM, which employs an RL agent with a) residual action and observation spaces based on affordances identified by FMs and b) LLM-guided exploration.
  2. 2.
    We introduce a prompting method for LLM-based exploration using hierarchical language-model programs, leading to faster convergence.
  3. 3.
    We show that ExploRLLM outperforms policies derived solely from LLMs and VLMs and generalizes to unseen scenarios, tasks, and real-world settings without additional training.
@@ II-A Foundation Models for Planning in Robotics
Researchers have shown that LLMs can exhibit reasoning capabilities and generate plans in zero-shot or few-shot settings [3, 10], which is crucial for high-level planning in robotics. These models facilitate task-level planning by integrating environmental groundings, such as affordance value scores [11] or feedback [12], with their language groundings. Furthermore, LLMs can generate robot-centric code programs as representations for both task-level [13] and skill-level planning [14]. Additionally, VLMs are increasingly integrated into robotics as a perception module of environmental context. The integration of knowledge from LLMs and VLMs can facilitate the creation of perception-planning pipelines [4] and the construction of 3D value maps for zero-shot planning frameworks [5]. However, due to real-world uncertainty, directly applying VLMs and LLMs to zero-shot tasks may not guarantee success or safety. Therefore, in our research, we treat these actions as exploratory behaviors within an RL framework.
@@ II-B Foundation Models and Reinforcement Learning
Incorporating FMs into RL frameworks has notably improved RL’s effectiveness. In [15], the authors have implemented LLMs as proxy reward functions, demonstrating their utility in RL. In the context of RL for robotics, LLMs are also capable of generating reward signals for robot actions by connecting commonsense reasoning with low-level actions [16], self-refinement [17] and evolutionary optimization over reward code to enable complex tasks such as dexterous manipulation [18]. Regarding exploration, authors in [19] reward RL agents toward human-meaningful intermediate behaviors by prompting an LLM. LLMs are also utilized as an intrinsic reward generator to guide exploration for long horizon manipulation tasks [20]. Contrary to these studies, our approach employs LLM-generated code policies as exploratory actions rather than focusing on reward shaping. Simultaneously with our study, [21] introduced a method for improving the sample efficiency of reinforcement learning with LLM-generated rule-based controllers. In [21], the RL policy is regularized towards replay data generated with the LLM policies. Our method instead uses LLM-generated policies for exploratory actions and does not promote the RL agent to be close to the LLM-generated policies.
@@ III Problem Formulation
In this study, we focus on language-conditioned tabletop manipulation tasks and a detailed overview of the method is shown in Figure 2. Each manipulation task begins at timestep $t=0$ with a linguistically described goal, denoted by $\bm{l}_{t}$. The agent receives an observation $\bm{o}_{t}$, consisting of an overhead RGB-D image and the state of the end-effector. Similar to existing methods (e.g., Transporter [22]), the action space involves a pick and a place primitive, denoted as $\{\mathcal{P}_{\mathrm{pick}},\mathcal{P}_{\mathrm{place}}\}$, with each action parameterized by pick and place positions in a top-down view. We simplify this to a single motion primitive—either pick or place. This simplification makes the RL problem more tractable by eliminating the need to learn a feature representation for each primitive individually. The pick or place action is defined as a tuple containing the primitive index $k$ (00 for pick, $1$ for place) and a top-down view position, expressed as $\bm{x}$, i.e., $\bm{a}_{t}=(k_{t},\bm{x}_{t})$. At each time step, the agent receives a reward $r_{t}$ consisting of a dense reward component $r^{d}_{t}$ and a sparse reward $r^{s}_{t}$.
@@ IV-A Observation and Action Spaces
Our method leverages the strengths of LLMs and VLMs to reduce the observation space used for the RL framework. First of all, the LLM reformulates user-provided language commands into predefined templates and highlights the objects within these templates to form an interpreted command vector $\tilde{\bm{l}}_{t}$. An example is shown in Figure 2a, where “Put the green block in the blue bowl” is interpreted into the template “Pick the [pick_object] and place it in the [place_object]”. It is important to note that, within a given task setting, the number and category of objects do not change. Utilizing VLMs as open-vocabulary object detectors, our system identifies and encloses objects relevant to the task within bounding boxes from the image, represented by their locations $\bm{X}_{t}=[\bm{x}^{0}_{t},\bm{x}^{1}_{t},...]$. RGB-D visual inputs are segmented into crops based on bounding box positions, denoted as $\bm{M}_{t}=[\bm{m}^{0}_{t},\bm{m}^{1}_{t},...]$. This method improves the system’s robustness to detection-inaccuracies and varying object shapes. The interpreted commands $\tilde{\bm{l}}_{t}$, the positional data $\bm{X}_{t}$ and the image patches $\bm{M}_{t}$ are then integrated into the reformulated RL observation $\bm{s}_{t}$ together with the robot gripper state (open/closed).
As the VLM already extracts each object’s position $\bm{x}^{i}_{t}$, the action space is converted into an object-centric residual action space (see Figure 2b). The reformulated action space consists of a primitive index $k$, an object index $i$ and a residual position $\bm{x}^{\mathrm{r}}$, expressed as $\tilde{\bm{a}}_{t}=(k_{t},i_{t},\bm{x}^{\mathrm{r}}_{t})$. This residual position is then added to the position of the object $i$, i.e., $\bm{x}_{t}=\bm{x}^{i}_{t}+\bm{x}^{\mathrm{r}}_{t}$. This residual action allows the agent to pick or place objects at specific locations. This is, for example, needed when picking the letter O, and $\bm{x}^{i}_{t}$ denotes the center of the bounding box. In this case, the residual action $\bm{x}^{\mathrm{r}}_{t}$ is needed to prevent picking the letter O at its empty center.
@@
1
Input: state $\bm{s}_{t}$, high-level LLM policy $\pi_{\mathrm{LLM}}^{H}$, low-level LLM policy $\pi_{\mathrm{LLM}}^{L}$, RL policy $\pi_{\mathrm{RL}}$
Output: action $\tilde{\bm{a}}_{t}$
2 Parameter: threshold $\epsilon$
$j\sim U_{[0,1)}$
  // Uniform sampling
3
4if $j\leq\epsilon$ then
       $\bm{a}^{H}_{t}=(k_{t},i_{t})\leftarrow\pi_{\mathrm{LLM}}^{H}(\bm{s}_{t})$
        // High-level
       $\bm{x}^{\mathrm{r}}_{t}\leftarrow\pi_{\mathrm{LLM}}^{L}(\bm{s}_{t},\bm{a}^{H}_{t})$
        // Low-level
5       $\tilde{\bm{a}}_{t}=(k_{t},i_{t},\bm{x}^{\mathrm{r}}_{t})$
6 else
       $\tilde{\bm{a}}_{t}\leftarrow\pi_{\mathrm{RL}}(\bm{s}_{t})$
        // RL policy
7
return $\tilde{\bm{a}}_{t}$
Algorithm 1 Exploration strategy $\pi_{\mathrm{EXP}}$
@@ IV-B LLM-Based Exploration
Traditional deep RL algorithms (e.g., SAC [23], PPO [24]) do not inherently promote frequent visits to high-value states in high-dimensional state-action spaces, making vision-based tabletop manipulation tasks particularly challenging. In such cases, RL agents may struggle when successful outcomes are rare. Leveraging the planning capabilities of LLMs and the perception strengths of VLMs can help guide the exploration process more effectively by tapping into the rich prior knowledge within these FMs. The LLM-based exploration strategy, denoted as $\pi_{\mathrm{EXP}}$ in Algorithm 1, draws inspiration from the $\epsilon$-greedy strategy. Specifically, during the rollout collection at each timestep, the off-policy RL agent employs the LLM-based exploration technique if a sampled random variable falls below the threshold $\epsilon$. Otherwise, the action is selected according to the current RL agent’s policy, $\pi_{\mathrm{RL}}$, as detailed in Algorithm 1.
Inspired by Code-as-Policy (CaP) [14], our method employs the LLM to generate hierarchical language model programs, which are executed during the training phase as exploratory actions. The hierarchical language model programs include high-level $\pi_{\mathrm{LLM}}^{H}$ and low-level $\pi_{\mathrm{LLM}}^{L}$ policy code programs. A high-level plan primarily involves selecting robot action primitives and the objects to interact with based on the current state of the robot and the objects.
In contrast to high-level tasks, instructing low-level actions poses a more significant challenge because high-level states and actions are more accessible and can be represented as language. When dealing with low-level actions, the complexity of the state becomes considerably more intricate, particularly for image-based problems. Therefore, instead of a deterministic code policy, we instruct the LLM to produce a code policy $\pi_{\mathrm{LLM}}^{L}$ for generating an affordance map according to the input image. The low-level exploration behavior is derived from a stochastic policy that relies on the values within this affordance map. Although the code generated by LLMs lacks guaranteed feasibility and accuracy in robot environments, these models can generate potentially useful policy candidates, with the one exhibiting the highest success rate being selected as shown in Figure 3.
@@ V Implementation
V-A RL Agent
We use the Soft Actor-Critic (SAC) algorithm with modifications in the collecting rollout phase, detailed in Algorithm 1. Other implementation aspects remain consistent with the standard SAC approach in stable-baselines3 [25]. We employ two convolutional layers to transform every image patch into a vector $\bm{\phi}\in\mathbb{R}^{n\times d}$, where n is the number of objects captured by VLM and $d$ the dimension of each patch as encoded by the CNN. The vector is subsequently concatenated with the position, robot gripper state, and the extracted episodic language goal $\bm{\tilde{l}}$ to form a new vector $\bm{\phi}^{\prime}\in\mathbb{R}^{n\times d^{\prime}}$, where $d^{\prime}$ denotes the dimension of each patch’s vector following encoding and concatenation. It then goes to a self-attention layer. The output features from this layer then go into a two-layer MLP. The structure mentioned above is consistently utilized across all actor and critic networks.
V-B VLM Detection
Utilizing an open-vocabulary object detector ViLD [26], objects in the environment can be identified by given specific labels. However, implementing this model online during training is time-consuming, so ViLD is utilized solely in the evaluation phase. In the training phase, the ground truth in the simulation is used to determine the center positions of the bounding boxes. It is important to note that ViLD’s position detection in real-world scenarios is not always flawless. To simulate this imperfection, Gaussian noise with a standard deviation equal to half the radius of the image crop is applied to the ground truth positions.
V-C LLM Code Policy Generation